# Kasner Eons

### v1

@@ Abstract
As shown in the classic work of Belinski, Khalatnikov and Lifshitz, the approach to a generic space-like singularity in general relativity consists of a sequence of eras and epochs in which the metric is locally Kasner, connected by brief transitions. When quantum gravity effects are included, we argue that a new type of transition arises, giving rise to Kasner eons: periods which are dominated by emergent physics at each energy scale. We comment on the different ways in which the Einsteinian eon may come to an end and show explicitly how additional Kasner eons arise in the interior of a black hole solution due to higher-derivative corrections.
Essay written for the Gravity Research Foundation 2024 Awards for Essays on Gravitation
@@ II Terminating BKL: three scenarios
Kasner eons. When the comoving volume of the spatial slices becomes smaller than a certain threshold, the effects of higher-derivative terms kick in. This typically takes place at a proper time $\tau\sim\ell$ away from the singularity, where $\ell$ is the length scale of new physics. Assuming one can still make sense of a classical evolution beyond this time, one would find that the relations for the Kasner exponents (2) will suffer a sharp transition to a new set of constraints. This marks the transition from the Einstein gravity eon to a higher-derivative eon, where the dynamical evolution is drastically altered. In this new eon we also expect to have an analogous BKL picture: an ultra-local singularity possibly with Kasner eras and epochs governing the approach, but with modified Kasner transition rules.
@@ III A toy model
[…]
@@ IV Discussion
We have explored three possible scenarios for the post-BKL phase: a (possibly never-ending) tower of higher-curvature Kasner eons characterized by modified Kasner exponents and transition rules; the emergence of a finite-volume curvature singularity; and the appearance of inner horizons which may ultimately result in a null singularity. This list is not intended to be completely exhaustive. Additional possibilities — such as bounces — may also occur. However, here we have chosen to focus on those possibilities that are manifest already in the simplest cases.

### v2

# Kasner Epochs, Eras and Eons

@@ Abstract
As shown in the classic work of Belinski, Khalatnikov and Lifshitz, the approach to a generic space-like singularity in general relativity consists of a sequence of epochs and eras in which the metric is locally Kasner, connected by brief transitions. When quantum gravity effects are included, we argue that a new type of transition arises, giving rise to Kasner eons: periods which are dominated by emergent physics at each energy scale. We comment on the different ways in which the Einsteinian eon may come to an end and show explicitly how additional Kasner eons arise in the interior of a black hole solution due to higher-derivative corrections.
@@ II Terminating BKL: three scenarios
Kasner eons. When the comoving volume of spatial slices becomes smaller than a certain threshold, the effects of higher-derivative terms kick in, which follows from the fact that such terms scale as inverse powers of the volume. This typically takes place at a proper time $\tau\sim\ell$ away from the singularity, where $\ell$ is the length scale of new physics. Assuming one can still make sense of a classical evolution beyond this time, one would find that the relations for the Kasner exponents (2) will suffer a sharp transition to a new set of constraints. This marks the transition from the Einstein gravity eon to a higher-derivative eon, where the dynamical evolution is drastically altered. In this new eon we also expect to have an analogous BKL picture: an ultra-local singularity possibly with Kasner eras and epochs governing the approach, but with modified Kasner transition rules.
@@ III A toy model
is the Gauss-Bonnet (GB) density, $\lambda$ is a coupling constant with dimensions of length2 and $G$ is the Newton constant, which in five dimensions has dimensions of length3.
@@ IV Discussion
We have explored three possible scenarios for the phase arising after the termination of the Einsteinian eon: a (possibly never-ending) tower of higher-curvature Kasner eons characterized by modified Kasner exponents and transition rules; the emergence of a finite-volume curvature singularity; and the appearance of inner horizons which may ultimately result in a null singularity. This list is not intended to be completely exhaustive. Additional possibilities — such as bounces — may also occur. However, here we have chosen to focus on those possibilities that are manifest already in the simplest cases.